OpenStudy (anonymous):

(x^2 - 16)(x - 3)^2 + 9x^2 = 0 Hint: Transform the equation and substitute a secondary variable (which equals an expression of x)

OpenStudy (anonymous):

The expression expanded and simplified is:\[x^4-6 x^3+2 x^2+96 x-144=0\]and can be factored.\[\left(x^2-8 x+24\right) \left(x^2+2 x-6\right)=0 \]

OpenStudy (anonymous):

kk thx

OpenStudy (anonymous):

(x^2-16)(x-3)(x-3)+9x^2=0 (x^2-16)((x^2)-3x-3x+9)+9x^2=0 (x^2-16)((x^2)-9x+9)+9x^2=0 (x^4)-9(x^3)+2(x^2)+144x+144=0

OpenStudy (anonymous):

thx guys